OpenStudy (anonymous):

help the ratio of men to women in a certain factory is 3 to 4. there are 216 men. how many workers are there?

OpenStudy (texaschic101):

3/4 = 216/x cross multiply (3)(x) = (4)(216) 3x = 864 x = 864/3 x = 288 check... 3/4 = 216/288 cross multiply (3)(288) = (4)(216) 864 = 864 (correct) we now know that there are 216 guys and 288 girls. In total, there are 504 workers

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

yesssss saidhari bought a toaster oven for $105. this list price was 1/4 off the list price. what was the list price?

OpenStudy (texaschic101):

the list price is 1/4 off the list price ??

OpenStudy (anonymous):

this price was 1/4 off the list price @texaschic101

OpenStudy (texaschic101):

1/4x = 105 -- divide by 1/4 x = 105/(1/4) x = 105 * 4/1 x = 420 check... 1/4(420) = 105 420/4 = 105 105 = 105 (correct) Regular price is $420. 1/4 off of 420 = 105

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

sure thing :)